HELP ME SOLVE THESE I'M TRYING TO TEACH MYSELF ALGEBRA FOR TO RETAKE THE COMPASS TEST, THANKS (a)8+(-5) (b)(-17)+(-12) (C) -9+7 (d) -9-(-12) (e) 23-(-13) (f) -18-(32) (g) (-7)(-12) (h) (-12)(8) (i) (8)(-9) (-48)+6 IF U CAN SHOW ME SOME FRACTION THAT WILL BE NICE, THANKS
Answer by Alan3354(69443)

(a)8+(-5) = 8-5 = 3
---------------------
(b)(-17)+(-12) = -17 -12 = -29
-----------------------
(C) -9+7 = -2
---------------
(d) -9-(-12) = -9 + 12 = 3
---------------
(e) 23-(-13) = 23 + 13 = 36
-----------------
(f) -18-(32) = -50
--------------
(g) (-7)(-12) = 84 (if they're multiplied)
---------------
(h) (-12)(8) = -96
---------------------
(i) (8)(-9) = -72
----------------
(-48)+6 = -42